What is the volume of 0.98 mil oxygen (02) gas at 275k and a pressure of 2.0 atm
1
Expert's answer
2019-09-12T03:32:40-0400

Volume of 0.98 mol of oxygen at S.T.P.= "0.98\times 22.4\ l=21.952\ l"

Now

Let "P_1,V_1,T_1\ be \ the \ quantities\ at \ S.T.P."

"V_1=21.952\\P_1=1\ atm\\T_1=273\ K\\P_2=2\ atm\\T_2=275\ K\\Also,

\frac{P_1V_1}{T_1}=\frac{P_2V_2}{T_2}"

"V_2=\frac{P_1V_1T_2}{P_2T_1}=\frac{1\times 21.952\times 275}{2\times 273}=11.056\ l"

Volume occupied by oxygen will be "11.056\ l"